OpenStudy (anonymous):

Solve the exponential equation using the method of relating the bases by first rewriting the equation b^u=b^v 6^x2/36^x=216

OpenStudy (amistre64):

well, 6^3 = 216, i remember that from a test I took a long time ago for some reason

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

so (-1,3)?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

6^(x^2) ------- = 6^3 ; unless I see it wrong, this is saying that 1 = 216 (6^2)^x)

OpenStudy (anonymous):

crud, it says (-1,3) isn't right

OpenStudy (anonymous):

really? 6^(x^2)/(36^x)=216 => 6^(x^2-2x)=216 => x = -1 or x = 3. is the problem written correctly?

OpenStudy (anonymous):

\[6^{x2} \div36^x=216\]

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=6^%28x^2%29%2F36^x+-+216 tehm is correct, but if your working with a program its prolly wanting a specific format

OpenStudy (anonymous):

thanks, I formatted the answer wrong. doh! You had the right one.

OpenStudy (amistre64):

:) i loathe working with programs ....